These are the news starting 01 Jan 2001; the older ones have been put to file
|16 Nov 2001||CO2 sensor repaired (membrane of pump was bad)|
|14 Nov 2001||CO2 sensor malfunction..Please be patient!|
|13 Nov 2001||Name change of the server: we are now http://meteo.lcd.lu. All old links are still functional, but it would be a good idea to update the bookmark!|
|23 Oct 2001||MeteoLCD will be unavailbale today from 10:00 UTC to about 16:00 due to a complete mains power shutdown .|
|03 Oct 2001||Added computation
and graph of windchill to the plot of windspeed; the windchill
is computed using the new
formula introduced this fall in Canada and the USA. The
computation is meaningless for windspeeds < 1.4 m/s (~5 km/h), so these
points will not be plotted. To give a better view, windchill and air
temperature are plotted together against the left y-axis.
For higher temperatures, the computed windchill may be higher than the actual air temperature!
There might be some initial hickups, so beware! A more thorough explanation on the new windchill computation can be found here...
|01 Oct 2001||Maintenance work and recalibration of the NOx sensor at 07:00. This explains the NO peak at that time.|
|16 Sep 2001||The station has worked nearly inattended from 10thAug. on (date of the last calibration); calibration and maintenance work was done the 13th Sep., with a complete download of the logger and a restart. Jul01 and Aug01 data files uploaded.|
|22 Jun 2001||Recalibration of ozone, NOx and CO2 sensors (this is the reason for the CO2 peak!)|
|19 Jun 2001||As always when touching a running system, there were a few hiccups: today the cumulative rain displayed on the graphs was the globe temperature. This has been corrected (19:32 UTC)|
|18 Jun 2001||Added a Babuc dry bulb sensor (PT100, logger channel 22), which will be used for the WBGT instead of the air tempearture sensor used until now. The graphs will be somewhat corrupted during the next days, please be patient!|
|29 May 2001||Pump of NOx sensor has been repaired at 06:30; sensor is operational again, and will be recalibrated asap|
|27 May 2001||The membrane of the vacuum pump of the NOx sensor became torn two during the 25th May, so that a good low pressure is not available anymore. The sensor was shut down yesterday, and will repaired asap. The main page (today) with the graphs will display a missing symbol in place of the NOx icon. .|
|18 May 2001||Maintenance work and recalibration on the CO2, NOx and O3 sensors; also checks on the WBGT. This work explains the "abnormal" spikes, which will be edited in the final data files.|
|13 May 2001||We had a bad problem starting 12May01 14:00: Due to an unknown cause (power loss? extreme temperature in the container = 36°C ???) the Linux machine showed a hard-disk problem. This has been corrected today at 07:00. All computer cases have been opened to help avoid extreme internal temperatures; we will install an UPS tomorrow to avoid the possible power-loss.|
|04 May 2001||We are still testing from time to time in-situ the calibration of the WBGT sensors, by disconnecting them for a short time from the logger and driving them by a special reader. This explains the "mysterious" wetbulb peak of this afternoon. This test-phase will continue for some time, so please be patient....|
|02 May 2001||Installed
Black Globe and Wet Bulb thermometers to measure the WBGT (Wet Bulb Globe
Temperature) which is a heat stress index. See picture!
Please disregard the PREVIOUS graph: only the last part starting 02May01 is ok
|29 Apr 2001||A revised version of the intercomparison report on the total ozone data from Microtops II at Diekirch and Brewer at Uccle is here!|
|27 Apr 2001||Everything is working again; we still have a problem with the automatic scheduler. We will be again on the air probably tomorrow evening...|
|26 Apr 2001||The day of the great lift: see the album for the pictures. At 19:00 we ahave reinstalled the major part of the equipment, but much electrical work remains to be done.|
|24 Apr 2001||We are moving again...The
MeteoLCD container will be displaced to another location on the roof of
our building. This needs a big crane and much work to dismount and remount
all equipment. Let's pray the Lord that nothing will break! I shall try to
take some pictures of this difficult procedure. We plan to be online
before next Sunday 29 April. This will (hopefully!) be the last move
before the completion of our new facilities, scheduled to be available at
the end of the year.
See the album of our "roof-movies"!
|14 Apr 2001||New version of the README.html file describing all sensors of the station|
|30 Mar 2001||1. Maintenance work on the NOx sensor: this
explains high NO peak
2. There are double (or sometimes missing lines) in the ASCII files today.dat, 7days.dat and all.dat. (the script for updating all.dat and correcting the other files need some tinkering...). This has NO influence on the monthly data files (which are fetched directly from the logger) and no visible influence on the graphs.
|21 Mar 2001||1. Recalibrated CO2 and NOx sensors
2. Added new script to compute cumulative rainfall (today and previous days)
|15 Mar 2001||UVB_Biometer problem repaired, probably due do a bad ground wire contact. The peak at 08:00 is caused by testing... please disregard and do not look for sudden sun supernova!|
|13 Mar 2001||There is a problem with the UVB Biometer, probably a bad electrical contact. Will be repaired asap...|
|10 Mar 2001||1. We had an interruption between logger and
one computer for several hours; things are ok now.
2. The monthly files jan01.dat and feb01.dat will be uploaded very soon; sorry for the delays caused by enormous workload... with coming springtime, everything will work better!
3. New CO2 Gasgard sensor is operational and works fine; has been fully recalibrated.
4. The small beautified icons are finally ok ( I use mogrify from ImageMagick to do the job)
|08 Mar 2001||1. NOx sensor repaired and working again
(starting 09:00 UTC)
2. Still some cosmetic problems with the thumbnails...
3. A new Gascard II CO2 sensor from Edinburgh Sensors will be installed tomorrow
|06 Mar 2001||1. Severe malfunction of the NOx sensor (the chopper motor
is out of order); no NO2 and NO data for the moment.
2. The CO2 sensor has been replaced by an old model, which works by diffusion. The graphs should display the correct concentrations; a calibration function must applied to the stored data: [CO2] = 1.261*data + 145.02. This function is applied to the graphed data!
|Still big problems with the CO2 sensor;
recalibration was not very successfull!
The somehow "pale" thumbnails will be bettered tomorrow
|03 Mar 2001||1. New modified interface for the TODAY data page, with
thumbnails and new sensor data added (more ASAP)
2. Rectified UVI computation (there was an error in a conversion factor)
3. Recalibrated CO2 sensor
|23 Feb 2001||Big calibration problems with the CO2 sensor...
be patient, I can cannot do anything before end of next week!
You can find here a BIG (3.5.MB) Excel file containing the full edited data set from 1998 to 2000, with averages, availability, total solar energy etc....
|22 Feb 2001||Some minor gliches and some fiddling on the new system
caused a temporary malfunction... should be ok now. The file all.dat
(huuuge file!) contains the gaps corresponding to missing data. The
monthly files (will) have all data.
The exhaust of the operators breathing is responsible for the high CO2 peak: it is not a sign of an impeding CO2 catastrophe, and it will NOT cause any global warming :-)
|19 Feb 2001||We started the new Linux system today at 15:00. Graphs are now in open-source, license-free PNG format, and will be updated every half hour. Thanks to M. Jean Mootz for help! (you may click here for two pictures taken today...). A detailed description of the new software handling will be written ASAP.|
|13 Feb 2001||Maintenance work on O3, NOx and CO2 sensors (shutdown of CO2 sensors explains negative peak...)|
|25 Jan 2001||Logger restarted, because file lcd_uv.dat became too long. 7 days graph will be rebuilt during the next days.|
|Back to Home-Page of Meteo_LCD|